# PQA: Zero-shot Protein Question Answering for Free-form Scientific Enquiry with Large Language Models

### v1

@@ 5.2 Zero-shot PQA:
PQA Learning with Novel Proteins:
5.3 Ablation Studies
To identify key contributing factors, we compared model performance across various dimensions. Most notably, reducing the size of the LLM from Phi-2, which has 2.8 billion parameters, to GPT-2 Medium, with 355 million parameters, resulted in a 20% decrease in the key “exact cofactor” metric (Table 7). However, the size of the ESM2 model did not seem to have a significant impact, with performance remaining largely similar between ESM2-S, ESM2-M, and ESM2-L (8M, 65M and 350M parameters, respectively). This observation could be attributed to the richness of representations in the Pre-trained PLMs, potentially highlighting a bottleneck in the LLM or the quality and quantity of the data (Table 7). Concordantly, the comparison of results from the zero-shot experiment also confirms that with a lower volume of data, a tailor-made dataset could yield better results. However, as the dataset size increases—a scenario analogous to the image domain—this factor becomes less crucial. These findings suggest that while LLM capacity significantly influences PQA model performance, the sophistication of sequence embeddings provided by ESMs reaches a point of diminishing returns, underscoring the importance of focusing on LLM enhancements and data quality for future improvements.
6 Conclusion
We’ve established a pioneering framework for zero-shot PQA, presenting a significant step forward in the application of LLMs for scientific enquiry. The introduction of specialized datasets and biologically relevant benchmarks underpins future explorations at the intersection of computational biology and artificial intelligence. Key insights from our zero-shot evaluations and ablation studies highlight the crucial role of LLMs in multimodal PQA performance, underscoring the importance of the knowledge depth and linguistic capabilities of LLMs. In essence, this work not only sets a new standard for scientific PQA but also opens the door for the research community to further refine and expand upon our methodologies. Looking ahead, leveraging larger, more diverse LLMs may offer substantial gains in model generalization, driving us towards the goal of automating accurate scientific enquiry.
7 Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning and Biology. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
Acknowledgements
We would like to thank Dr Aida Nematzadeh for critical advice and Maryam Karaminejad Ranjbar for graphic design.
@@ Appendix A Supplementary Methods
A.1 Data collection from UniProt
@@
A.2 Debiasing SwissProt Entries Using UniRef50 Clusters
@@
A.3 Example Instructions to GPT3.5
You will receive details about a specific protein. Perform the following tasks: 1) Provide a factual summary, with a maximum of 500 words, that accurately and scientifically describes the functional, biochemical and structural properties of this protein based only on the provided information. Ensure that the summary follows a natural and scientific flow, starting with general information such as structure, localization and taxonomy before detailing functional and biochemical properties. Ensure that all key points are covered and DON’T provide any further information than what is stated in the input. 2) For each type of information provided, create a question-and-answer pair to elucidate an aspect of the protein’s functionality or biochemical properties without using the protein’s name.
Use the below ; separated two-element tuples structure for the output: (summary: this protein…); (what is question 1?, the answer 1); (what is question 2?, the answer 2); …
For both tasks if the input contains large group of properties only provide the canonical and crucial information rather than enumerating every single entry. DON’T use any of your knowledge to add additional context or information. DON’T add any speculative or unwarranted information that is not specifically provided. AVOID using generic phrases or embellished terms such as ’highly’, ’several’, ’diverse’ and ’various’.
A.4 Biochem-Lite Metric Questions and Motivation
@@
A.5 MLP Baseline
We use a 3-layer MLP with GELU activation on protein sequence embeddings from the Pre-trained PLM. Where possible, we convert each light-weight benchmarking question into a classification or value prediction task (Table 2). The MLP is then trained to predict the correct class labels or values using cross entropy or mean squared error loss, respectively. This baseline sets an upper limit for the information content inherent in PLM embeddings, with any improvements potentially indicating generalization due to the knowledge encapsulated in the LLM.
A.6 Hyperparameter Optimization
@@
Appendix B Supplementary Figures and Tables
@@
Figure B.1: Over-representation bias in SwissProt database.
Figure B.3: Evaluating the Effectiveness of Biochem-Lite Versus Traditional Linguistic Metrics for Scientific Accuracy of PQA. Statistical significance is determined through a one-tailed paired t-test across three randomly seeded data subsets and model training (significance guide: ${}^{*}p<0.05$, ${}^{**}p<0.01$, ${}^{***}p<0.001$). By design, Self-Pika is expected to outperform Pika w/o PLM across all metrics, except for the mw MALE metric.

### v2

@@ 4.2.6 Ablation Studies
5 Conclusion & Impact
We’ve established a pioneering framework for zero-shot PQA, presenting a significant step forward in the application of LLMs for scientific enquiry. The introduction of our specialized datasets and biologically relevant benchmarks underpins future explorations at the intersection of computational biology and artificial intelligence. Key insights from our zero-shot evaluations and ablation studies highlight the crucial role of LLMs in multimodal PQA performance, while underscoring the importance of considering evolutionary relation in assessing performance. Since, for Pika-DS we employed GPT3.5 to create synthetic annotations, given the limitations of LLMs in generating large-scale synthetic datasets, we endeavored to minimize the inclusion of harmful content in Pika-DS through prompt optimization and manual evaluation. Nonetheless, due to the dataset’s extensive size, there’s a slight possibility that unintended harmful content might still be present. Our Pika-based pre-trained models are derived from the publicly accessible and unmoderated Phi-2 LLM. Thus, all cautions, restrictions, and notes associated with Phi-2 (Microsoft, 2023) are applicable to our models. Looking ahead, leveraging larger, more diverse LLMs may offer substantial gains in model generalization, driving us towards the goal of automating accurate scientific enquiry into proteins.
6 Reproducibility & Accessibility
All code and data used for creation of data, model training and baselines are publicly accessible on www.github.com/EMCarrami/Pika.
The Pika framework is specifically designed for question answering related to protein sequences. With scientists having identified nearly 0.25 billion protein sequences, and functional annotations available for fewer than a million, our framework offers significant potential for research into these largely unexplored proteins. While our efforts are directed towards scientific research, we recognize the potential risk of misuse by individuals aiming to create or identify harmful substances. We strictly prohibit using our dataset and framework for any illegal activities or actions that could harm individuals or society.
Acknowledgments and Disclosure of Funding
We would like to thank Dr Aida Nematzadeh for critical advice and Maryam Karaminejad Ranjbar for graphic design. We would also like to sincerely thank Relation Therapeutics for providing us with the necessary compute for training the final version of Pika. This work was primarily supported by personal funding and the authors declare no conflict of interest.
@@ Appendix A Supplementary Methods
A.1 Creation of Pika-DS
Pika-DS was prepared in three phases:
  1. 1.
  2. 2.
    UniProt also offers similarity-based clustering of its entries at various sequence similarity thresholds of 100%, 90%, and 50%, known as UniRef clusters. For instance, protein sequences that share a minimum of 50% pairwise sequence similarity will be assigned the same UniRef50 cluster. Recognizing a strong bias towards more commonly studied protein families in SwissProt (Figure B.1), we limited Pika-DS to a maximum of two most informative sequences per UniRef50 cluster using a custom algorithm (Section A.3). This resulted in a debiased set of over 257,000 protein sequences and 1.17 million information fields.
  3. 3.
    Lastly, we used GPT3.5 API to process the information fields for each protein entry using systematically optimized prompts (Sections A.4 and A.5) to create the Pika-DS’s three main components (example in Table B.2):
    • QAs: Several diverse QA pairs for each information field, formatted for LLM training.
    • Metrics: Single-word answers to a set of predefined scientific questions serving as the ground-truth for our Biochem-Lite benchmarks (Section 3.2.2).
A.2 Data collection from UniProt
@@
A.3 Debiasing SwissProt Entries Using UniRef50 Clusters
@@
A.4 GPT3.5 Prompt Optimization Strategy
To ensure the high quality of GPT3.5 generated information and QA pairs, we performed a systematic prompt optimization focusing on GPT3.5’s adherence to using only the given data, summarizing complex biological reactions accurately, and avoiding speculative or ambiguous language. This was performed in an iterative process where at each iteration, GPT3.5 was provided with the prompt and all the extracted information for 50 randomly selected protein sequences. Next, 100 outputs from each of the summary statements, QA pairs and evaluation metrics were evaluated against the input information of the respective proteins by an expert biochemist and the number of incorrect, poor quality and irrelevant (correct but unrelated to the protein’s function) statements or QA pairs were noted. At each step, the prompt was modified to address the most problematic issues, with a target of no more than one incorrect, two poor quality and two irrelevant values in each category (optimised prompt in Section A.5).
A.5 GPT3.5 Prompts Used for Creation of Pika-DS
Summarising and QAs:
The following instructions were used to create summary statements and QA annotations based on information fields collected from SwissProt:
You will receive details about a specific protein. Perform the following tasks and print each result in a new line:
1) Provide a factual summary, without using the protein’s name with a maximum of 500 words. Your summary must accurately and scientifically describe the functional, biochemical and structural properties of this protein based only on the provided information. Ensure that the summary follows a natural and scientific flow, starting with general information such as structure, localization and taxonomy before detailing functional and biochemical properties. Ensure that all key points are covered and DON’T provide any extra information than what is stated in the input.
2) For each type of information provided, create a question-and-answer pair to elucidate an aspect of the protein’s functionality or biochemical properties without using the protein’s name. Phrase your questions and answers such that they will be suitable for training a language model. DON’T enumerate or label the questions and print each question and its answer pair in the same line.
- For all tasks if the input contains large group of properties only provide the canonical and crucial information rather than enumerating every single entry. - Where applicable, summarise enzymatic reactions into one or two of the generic classes of the activities. - DON’T use any of your knowledge to add additional context or information. DON’T add any speculative or unwarranted information that is not specifically provided. - AVOID using generic phrases or embellished terms such as ’highly’, ’several’, ’diverse’ and ’various’. - Exactly follow the output format provided below to ensure consistency. Final output format: summary: [YOUR SUMMARY] QA pairs: 1) [YOUR QUESTION] [YOUR ANSWER] 2) [YOUR QUESTION] [YOUR ANSWER] […] n) [YOUR QUESTION] [YOUR ANSWER]
Biochem-Lite Ground truth:
The following instructions were used to collect ground-truth values for Biochem-Lite questions based on summary statements and information fields:
You will receive details about a specific protein. Provide a single word answer to the following questions. Print each question and your answer in the same new line. If the question does not apply to the protein, ignore the question. 1) Is this protein localized to the cell membrane? 2) Is this a membrane protein? 3) Is this protein localized to nucleus? 4) Is this protein localized to mitochondria? 5) Does this protein bind to DNA? 6) Does this protein bind to RNA? 7) Is this protein an enzyme? 8) What are co-factors of this protein as a comma separated list?
- Exactly follow the output format provided below to ensure consistency. Final output format: 1) [MY QUESTION] [SINGLE WORD ANSWER e.g. YES/NO/UNKNOWN] 2) [MY QUESTION] [SINGLE WORD ANSWER e.g. YES/NO/UNKNOWN] […] 7) [MY QUESTION] [SINGLE WORD ANSWER e.g. YES/NO/UNKNOWN] 8) [MY QUESTION] [Comma separated list of co-factors e.g. FAD,FMN/UNKNOWN]
A.6 Biochem-Lite Metric Questions and Motivation
@@
A.7 Metric equations used in Biochem-Lite
  • mw MALE: The mean-absolute log error (MALE) of the predicted molecular weight (MW) is:
    $MALE=\frac{1}{N}\sum_{i=1}^{N}|\log_{10}(\hat{MW}_{i}/MW_{i})|$ (1)
  • exact cofactor: The score is computed as:
  • location F1: The F1 score is the harmonic mean of Precision and Recall. The predicted class assignments for the calculation of Precision and Recall are based on the exclusive presence of correct labels (membrane, nucleus, or mitochondrion) in the generated response. Responses lacking or containing multiple labels receive the none class label.
  • is_enzyme F1: The F1 score is computed similar to location F1. Responses are classified as True or False based on the exclusive presence of yes or no, respectively. Any deviation is labeled as none.
  • binary Q mean F1: We have:
    $F1_{\text{binary Q mean}}=\frac{1}{3}\sum_{q=1}^{3}F1_{q}$ (3)
    where $F1_{q}$ is the F1 score for each binary question (class assignments similar to is_enzyme F1).
A.8 MLP Baseline
We use a 3-layer MLP with GELU activation on protein sequence embeddings from the Pre-trained ESM2. Where possible, we convert each light-weight benchmarking question into a classification or value prediction task. More specifically we performed regression for log values of mw for mw MALE and performed classification for is_enzyme and location questions. The MLP is then trained to predict the value or correct class labels using mean squared error or cross entropy loss, respectively. This baseline sets an upper limit for the information content inherent in ESM2 embeddings.
A.9 Training
We devised a simple training strategy for PQA models, keeping both the PLM and LLM frozen and optimizing for the causal language model loss with AdamW. Hyperparameters were optimised following a greedy search as detailed in section A.10. All training were performed on a single A100-80GB or H100-80GB Nvidia GPU. Unless otherwise stated, both the QA and summary statement section of the Pika-DS was used for the training of all Pika models. Examples were split to train, validation and test set based on the UniRef50 cluster or EvoGroup of their respective protein sequences in 94.5%, 0.05% and 5% ratios, respectively.
A.10 Hyperparameter Optimization
@@
A.11 Creation of EvoGroups
EvoGroups were decided by starting from a random sequence in the data, identifying all its related sequences using JackHammer (Potter et al., 2018) with a very lenient cut-off threshold for e-E-score of 1.0 to ensure a broad grouping of related proteins. For each randmoly selected sequence, all related sequences were marked as belonging to the same EvoGroup and were removed from the remainder of the dataset for further iterations with newly selected sequences. The process was repeated until all sequences were assigned an EvoGroup.
Appendix B Supplementary Figures & Tables
Figure B.1: Over-representation bias in SwissProt database.
# total # unique # words # tokens
Proteins 257 K 252 K - 105.3 M
Info fields 1.17 M 319 K 10.8 M 29.5 M
Summaries 1.33 M 861 K 22.5 M 36.4 M
Questions 1.71 M 98 K 15.7 M 14.9 M
Answers 1.71 M 783 K 18.4 M 31.8 M
Table B.1: Pika-DS size and token counts. Values for Summaries represent the number of sentences.
@@ Appendix B Supplementary Figures & Tables
Random
LLM
only
self-Pika
w/o PLM
self-Pika
F1 Score 0.49 0.00 0.43 0.99
Accuracy 0.50 0.00 0.51 0.99
Table B.3: Performance on Control Question.